### v1

# Bio-Inspired Classification: Combining Information Theory and Spiking Neural Networks - Influence of the Learning Rules

Abstract
Training of Spiking Neural Networks (SNN) is challenging due to their unique properties, including temporal dynamics, non-differentiability of spike events, and sparse event-driven activations. In this paper, we widely consider the influence of the type of chosen learning algorithm, including bioinspired learning rules on the accuracy of classification. We proposed a bioinspired classifier based on the combination of SNN and Lempel-Ziv complexity (LZC). This approach synergizes the strengths of SNNs in temporal precision and biological realism with LZC’s structural complexity analysis, facilitating efficient and interpretable classification of spatiotemporal neural data. It turned out that the classic backpropagation algorithm achieves excellent classification accuracy, but at extremely high computational cost, which makes it impractical for real-time applications. Biologically inspired learning algorithms such as tempotron and Spikprop provide increased computational efficiency while maintaining competitive classification performance, making them suitable for time-sensitive tasks. The results obtained indicate that the selection of the most appropriate learning algorithm depends on the trade-off between classification accuracy and computational cost as well as application constraints.
Index Terms:
spiking neural networks, learning algorithms, Artificial Intelligence, machine learning, networks dynamics, Shannon Information Theory, Lempel-Ziv Complexity, signal processing
I Introduction
Backpropagation is a gradient-based learning rule commonly applied in networks training [1, 2]. Artificial Neural Networks (ANNs) and Spiking Neural Networks (SNNs) are both inspired by biological neural systems, however, they differ from the computation point of view. While ANNs dominate deep learning due to their simplicity, SNNs better mimic real neurons by using spike-based communication. However, training SNNs to match ANN accuracy, especially with backpropagation-based methods [1], requires significant hardware resources and is computationally expensive. As a consequence, SNNs training remains a challenge due to their complex dynamics and the non-differentiable nature of spike-based activity [3].
Learning rules can be categorized into three main types: unsupervised learning (for example, Hebian learning [4], Symmetric Spike-Driven Synaptic Plasticity (SDSP) [5, 6]), and Spike Timing Dependent Plasticity (STDP) [8, 7]), supervised learning (like backpropagation (BP)[1, 2], Spike-Timing-Dependent Backpropagation (STBP) [10, 9], tempotron [13, 12, 11], SpikeProp [14, 15], Chronotron [16, 17], and Remote Supervised Method (ReSuMe) [18]), and hybrid learning (including Artificial-to-Spiking Neural Network conversion (ANN-SNN conversion) [19], and Reward-based STDP [20]).
A commonly used training method in SNNs is SpikeProp, which is an arithmetic-based learning algorithm conceptually similar to the BP algorithm in ANNs [21]. However, arithmetic-based learning rules may be not a right choice for building biologically efficient networks. To achieve more biologically plausible learning, several alternative methods have been developed. For example, STDP that is inspired by biological synaptic plasticity, where neurons strengthen or weaken connections based on temporal contiguity. However, despite its efficiency in synaptic adaptation, STDP does not match the classification performance of BP-based ANN training [22, 23]. Moreover, spatio-temporal backpropagation (STBP) extends the traditional backpropagation algorithm to handle both spatial and temporal dependencies in spiking neural networks. This makes it possible to train deep SNNs using gradient-based optimization [24]. Other commonly used learning algorithms in SNN training are ReSuMe [18, 10, 25], and Chronotron [16, 17]. They are effective in adaptation to structured spike patterns, however, both are computationally demanding and require significant training time. An interesting SNN learning rule, namely tempotron was proposed by [13]. It is particularly effective for temporal coding, where information is embedded in the precise timing of spikes. In the study [26] training is conducted using the tempotron learning rule, while optimization is achieved through STDP. This combination significantly improves processing speed and reduces computational complexity compared to one can say "conventional" spiking neural network methods. In turn, [27, 28] proposed a neuron normalization technique and an explicitly iterative neuron model, which resulted in a significant increase in the SNNs’ learning rate. Another training strategy involves first training an ANN (typically using perceptrons) and then converting it into an SNN with an equivalent structure [19]. This conversion leverages the strengths of ANN training while allowing SNNs to benefit from energy-efficient spike-based processing. However, a major drawback is that this approach fails to capture temporal spike dynamics, as reliable frequency estimation requires a non-trivial time period. On the other hand, to optimize data efficiency, various active learning approaches have been introduced, including Bio-Inspired Active Learning (BAL), Bio-Inspired Active Learning based on Firing Rate (BAL-FR), and Bio-Inspired Active Learning utilizing Membrane Potential (BAL-M) [29]. These methods identify and select training samples by analyzing the discrepancies between empirical observations and generalization behavior, ensuring that previously unrecognized patterns are incorporated into the learning process. In turn, [9] introduces a Fractional-Order Spike-Timing-Dependent Gradient Descent (FO-STDGD) learning model, which applies fractional gradient descent principles to improve learning in SNNs.
We propose a novel hybrid approach that combines the temporal precision and biological plausibility of spiking neural networks with the Lempel-Ziv complexity (LZC) measure [41] to enhance classification performance. To our knowledge, this is the first time LZC has been integrated with SNNs. This method enables efficient and interpretable classification of spatiotemporal neural data by quantifying the structural complexity of spike patterns. It also improves robustness to noise and is particularly effective for data with variable temporal dynamics, such as Poisson-distributed signals. It turned out that the capability of bio-inspired neural networks in biosignal classification is highlighted, emphasizing their high precision, sensitivity, and reliability in distinguishing physiological patterns.
II Mathematical Formulation
II-A Notation
In Table I the basic notation is presented.
Notation Description
$\textbf{x}=[x_{1},x_{2},\cdots,x_{n}]\in\mathbb{R}^{n}$ synaptic weight vector
$V_{m}(t)\in\mathbb{R}$ membrane potential at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$I(t)\in\mathbb{R}$ input current at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$t_{f}\in\mathbb{R}^{+}$ firing time of neuron
$S_{i}(t)\in\mathbb{Z}$ binary-valued spike train (0 or 1 at any given time)
$t_{i}^{k}\in\mathbb{R}^{+}$ spike times (real-valued, positive)
$\Theta(V_{n})\in\{0,1\}$ Heaviside function (1 if neuron fires, 0 otherwise)
$\eta\in\mathbb{R}^{+}$ general learning rate (positive real number) for weight update in supervised learning and gradient-based optimization
$\eta_{+},\eta_{-}\in\mathbb{R}^{+}$ STDP time constant (positive real)
$A\in\mathbb{R}$ SDSP learning rate (real-valuates), defines the strength of the symmetric weight update rule
$A_{-},A_{+}\in\mathbb{R}^{+}$ amplitude of STDP weight changes (positive real)
$t_{pre},t_{post}\in\mathbb{R}^{+}$ pre- and postsynaptic spike time (positive real)
$t_{spike}\in\mathbb{R}^{+}$ spike timing
$S_{pre}(t),S_{post}(t)\in\{0,1\}$ binary spike trains
$K(t)\in\mathbb{R}\longrightarrow\mathbb{R}$ kernel function, maps time to real value
$t_{i}^{target},t_{i}^{actual}\in\mathbb{R}^{+}$ targeted and ctual spike times
$E\in\mathbb{R}^{+}$ error function (positive real)
$\frac{\partial E}{\partial w_{i}}\in\mathbb{R}$ gradient of error
$\Phi(t_{pre},t_{post})\in\mathbb{R}^{+}$ loss function (positive real)
$S_{teach}(t)\in\{0,1\}$ binary teacher spike train
$w_{SNN},w_{ANN}\in\mathbb{R}$ synaptic weight in SNN/ANN (real-valued)
$\tau_{syn}\in\mathbb{R}^{+}$ synaptic time constant (positive real)
$r(t)\in\mathbb{R}$ reward signal (real-valued)
$\Delta w_{STDP}\in\mathbb{R}$ STDP-based weight change (real-valued)
$U(w_{i})\in\mathbb{R}\longrightarrow\mathbb{R}$ uncertainty function that maps weight to real value)
TABLE I: Basic notation used in the study.
II-B Unsupervised Learning Algorithms
Unsupervised learning algorithms learn patterns and structures from unlabeled data without explicit supervision or target output, i.e., provide local synaptic modifications with eplicit error signals.
II-B1 Hebbian Learning Rule
Hebbian Learning Rule strengthens synapses when both neurons fire together [4]
$\Delta w_{i}=\eta(S_{pre}(t)S_{post}(t)).$ (1)
A major drawback of Hebbian learning is its decreasing efficiency as the number of hidden layers increases [30]. Although it remains competitive with up to four layers, its performance diminishes in deeper networks, limiting its scalability for complex architectures.
II-B2 Spike Timing Dependent Plasticity
STDP is a biologically inspired rule that adjusts synaptic weights based on the relative timing of pre- and postsynaptic spikes [8, 7]:
$\Delta w_{i}=\begin{cases}A_{+}e^{-\frac{(t_{post}-t_{pre})}{\tau_{+}}}&t_{post}>t_{pre}\\ -A_{-}e^{-\frac{(t_{post}-t_{pre})}{\tau_{-}}}&t_{pre}>t_{post}.\end{cases}$ (2)
It is a form of Hebbian learning (1). It is essential for modeling the temporal dynamics of biological neural systems. Unlike traditional learning methods, its event-driven nature enables unsupervised learning, allowing neural networks to adapt autonomously to temporal patterns in input data. This makes it valuable for processing spatio-temporal signals, such as sensory data. However, STDP requires careful parameter tuning, as factors like scaling and time constants significantly impact network behavior. In addition, its computational demands can be high, particularly in large-scale simulations [31].
II-B3 Symmetric Spike-Driven Synaptic Plasticity
SDSP is a variant of STDP, however, with symmetric weight updates [5, 6]
$\Delta w_{i}=A(S_{pre}(t)-S_{post}(t)).$ (3)
It is a biologically plausible learning mechanism that mirrors neural dynamics by focusing on individual spike events. This event-driven approach enables rapid adaptability, making it valuable for real-time learning. However, its sensitivity to single spikes poses challenges in noisy environments, requiring advanced filtering techniques [32, 33].
II-C Supervised Learning Algorithms
Supervised learning maps input data to output labels based on labeled dataset. It relies on error-based updates.
II-C1 Backpropagation Algorithm
Backpropagation is a gradient-based learning rule commonly applied in networks training [1, 2].
$\Delta w_{i}=-\eta\frac{\partial E}{\partial w_{i}}$ (4)
where $E$ is the error function. The weight update follows a chain rule
$\frac{\partial E}{\partial w_{i}}=\frac{\partial E}{\partial V_{m}}\cdot\frac{\partial V_{m}}{\partial w_{i}}.$ (5)
II-C2 Spike-Timing-Dependent Backpropagation
Spike-Timing-Dependent Backpropagation extends the conventional gradient-based learning rule of backpropagation to spiking neural networks, incorporating spike-timing-dependent plasticity principles to optimize synaptic weights [10, 9]. The weight update rule is given by Equation (4). In STBP, the gradient computation follows the chain rule:
$\frac{\partial E}{\partial w_{i}}=\frac{\partial E}{\partial t_{\text{spike}}}\cdot\frac{\partial t_{\text{spike}}}{\partial V_{m}}\cdot\frac{\partial V_{m}}{\partial w_{i}}.$ (6)
Unlike traditional backpropagation, STBP accounts for the temporal dynamics of spikes, leveraging precise spike timing to adjust synaptic strengths. This biologically inspired approach enables more efficient credit assignment in spiking networks while maintaining compatibility with gradient-based optimization methods.
II-C3 Tempotron Learning Rule
Tempotron learning rule adjust synaptics weight to classify temporal spike patterns [13, 12, 11]
$\Delta w_{i}=\eta\sum_{t\in T_{pre}}K(t_{f}-t)$ (7)
The tempotron learning principle is a biologically inspired approach that refines synaptic weights based on the precise timing of incoming spikes rather than just their frequency. Unlike conventional neural models that prioritize weight adjustments or network structures, the tempotron underlines the importance of timing, showing that the "when" of a neural event can be as crucial as the "where" or "how often".
II-C4 SpikeProp Algorithm
SpikeProp applies gradient descent to minimize spike timing errors [14, 15]:
$\frac{\partial E}{\partial w_{i}}=\sum_{t_{post}}\frac{\partial E}{\partial t_{i}^{actual}}\cdot\frac{\partial t_{i}^{actual}}{\partial w_{i}}.$ (8)
SpikeProp extends backpropagation to spike-based networks, enabling gradient-based optimization. It is computationally expensive due to temporal dynamics. Also, requires careful handling of non-differentiability in spike generation.
II-C5 Chronotron Learning Rule
The Chronotron minimizes the actual and target spike times [16, 17]
$E=\sum_{i}\sum_{k}\frac{(t_{i}^{actual}-t_{i}^{target})^{2}}{2}$ (9)
$\Delta w_{ij}=-\eta\sum_{k}(t_{i}^{actual}-t_{i}^{target})\frac{\partial t_{i}^{actual}}{\partial w_{ij}}$ (10)
A key advantage of this approach is its consistent coding for both inputs and outputs, improving the efficiency of SNNs. However, the granularity of the method can significantly increase computational demands, requiring precise parameter tuning and high-quality temporal data.
II-C6 Remote Supervised Method
ReSuMe combines STDP with supervised learning using a teacher spike train [18]
$\Delta w_{ij}=\eta\left[S_{pre}(t)*S_{teach}(t)-S_{pre}(t)*S_{post}(t)\right]$ (11)
ReSuMe enables supervised training while maintaining biological plausibility [34].
II-D Hybrid Learning Algorithms
Hybrid algorithms integrate elements of both supervised and unsupervised learning.
II-D1 Artificial-to-Spiking Neural Network conversion
ANN-SNN conversion maps pre-trained ANN weights to SNN synapses [19]
$w_{SNN}=\frac{w_{ANN}}{\tau_{syn}}.$ (12)
Converting ANNs to SNNs allows the use of existing training methodologies while leveraging SNNs’ energy efficiency and biological realism [35]. This conversion maps ANN activation values to the firing rates of SNN neurons, requiring precise threshold tuning to maintain accuracy. Improper thresholds can lead to excessive or insufficient spiking, affecting performance.
II-D2 Reward-based STDP
Reward-based STDP modifies STDP with reinforcement learning [20]
$\Delta w_{ij}=\eta r(t)\left[S_{pre}(t)*S_{post}(t)\right].$ (13)
II-D3 Bio-inspired Active Learning
Active learning updates weights based on uncertainty [37, 38, 36]
$\Delta w_{ij}=\eta\cdot U(w_{i})\cdot\mathbb{E}\left[I(S_{post};S_{pre})\right].$ (14)
Integrates biological learning mechanisms with active learning strategies to improve training efficiency [36]. Unlike traditional supervised learning, which relies on large labeled datasets, BAL actively selects the most informative data points for labeling, reducing data requirements while maintaining high performance.
III Classification Task
Figure 1: The architecture of the considered neural network.
The SNN neural network under consideration consists of three layers: input, hidden, and output, each containing $n$ neurons; see Figure 1. It was based on the Leakly-Integrate-and-Fire (LIF) biophysical neuron model that is a fundamental component in SNNs for simulating the subthreshold behavior of a neuron’s membrane potential $U(t)\in\mathbb{R}$ [40, 39]. Given an input vector $\mathbf{x}(t)\in\mathbb{R}^{n}$, synaptic weights $\mathbf{w}\in\mathbb{R}^{n}$, and bias $b\in\mathbb{R}$, the net input current is defined as:
$I(t)=\mathbf{w}^{T}\mathbf{x}(t)+b=z(t),$ (15)
where $z(t)$ denotes the weighted synaptic drive at time $t$. The membrane voltage evolves according to:
$\tau_{m}\frac{dU(t)}{dt}=-U(t)+z(t),$ (16)
and emits a spike when $U(t)\geq U_{\text{th}}$, after which the voltage is reset to a resting value $U_{r}$. This model captures the temporal integration and passive leakage of membrane charge in a computationally efficient form.
We considered different sets of neuron parameters and the presented results include the most efficient set of parameters for each learning algorithm. The threshold took values 1.00, 0.50, 0.40, 0.30, 0.20, 0.10, and 0.05. The decline was 0.100, 0.05, 0.03, 0.01, and the learning rates 0.0500, 0.0100, 0.0010, 0.0098, and 0.0001. We considered neural networks composed of 16, 32, 64, 128, 256, 512, and 1024 neurons in layers. The best of the obtained results are presented in Table II, Table III, and Table IV.
The network processes binary sequences of length 1024, which are transformed into $n$-bit spike trains and subsequently converted back into binary sequences. The classification of these output sequences is performed using the Lempel-Ziv complexity (LZC) measure, a well-established metric for evaluating the informational content of a sequence. A binary sequence $\textbf{x}_{n}^{1}:=[x_{1},x_{2},\dots,x_{n}]$, where each $x_{i}\in\{0,1\}$ the LZ complexity $C_{\alpha}(\textbf{x}_{n}^{1})$ represents the count of distinct substrings that appear sequentially in $\textbf{x}_{n}^{1}$. The normalized complexity, which assesses the rate at which novel patterns emerge, is defined as:
$c_{\alpha}(\mathbf{x}_{n}^{1})=\frac{C_{\alpha}(\mathbf{x}_{n}^{1})}{n}\log_{\alpha}{n},$ (17)
where $\alpha=2$ for binary sequences. For purely random sequences, $c_{\alpha}(\textbf{x}_{n}^{1})$ converges to 1, while for deterministic sequences, it approaches 0. This measure effectively approximates the entropy of ergodic stochastic processes, making it a valuable tool to evaluate randomness and structure within the output of the neural network [41, 42, 43]. For completeness, we performed our analysis for information sources for which the subsequent symbols are more and more correlated [44]. Thus, we consider three types of dataset that consist of Bernoulli sequences (uncorrelated), two-state Markov, and Poisson processes. The Poisson process has been often used to model spike trains since early studies, due to the proportionality between mean and variance in many neural responses. Experimental data have confirmed this Poisson property [45]. The binary sequence of each process is converted into inputs to LIF models with specified $\Delta t=1ms$ [11]. For the evaluation of the proposed classification task, we applied accuracy, the mean squared error (MSE), and the mean absolute error (MEA).
IV Results and Discussion
We consider classical and bio-inspired learning algorithms in SNNs. Table II presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. All computation were provided during 10 epochs, on Intel(R) Core(TM) i7-14700F, 2.10 GHz. A general finding of our study is the pronounced dependence of both classification accuracy and computational overhead on the choice of learning algorithm and the underlying characteristics of the input data. Observed classification accuracies ranged from 89.50% to 100.00% across the evaluated methods, whereas computational times exhibited substantial variation, ranging from approximately 7 seconds to over 48,000 seconds.
Table II presents the influence of the unsupervised learning algorithm such as Hebbian, STDP, and SDSP on the performance of SNNs. The results obtained for the Hebbian algorithm demonstrate that, within the classical framework, high classification accuracy was achieved for the Bernoulli and Markov processes, with values of 99.00% and 95.79%, respectively. These results were accompanied by relatively long computation times of 52.93 s and 74.66 s. Furthermore, MSE values of 0.0011 and 0.00422, along with MAE values of 0.0100 and 0.0422, respectively, suggest minimal prediction errors. The coefficient of determination $R^{2}$ scores of 0.9600 and 0.8313 further confirm a strong agreement between predicted and actual values. For the Poisson process, a classification accuracy of 90.00% was observed, with a $R^{2}$ score of 0.9074. The corresponding MSE and MAE values were 0.1000, while the $R^{2}$ score of 0.5998 indicates a moderate correlation between the predictions and the actual values. It is worth emphasizing that the accuracy for the Poisson process is achieved for a twice smaller number of neurons in the layers (for Bernoulli and Markov processes, we achieve similar accuracy with the same number of neurons in the layers). An increase in the number of neurons causes a decrease in accuracy below 50.00%, contrary to the case of Bernoulli and Markov processes, where accuracy increases with increasing number of neurons in layers. Thus, the worse classification results for the Poisson process may be related to the fact that events occur randomly in time, making it difficult to generalize Hebbian learning effectively. The relatively low computation time indicates that the Hebbian learning algorithm efficiently processes structured data. However, taking into account the nature of Hebbian learning it is effective when there are strong, predictable relationships between inputs and outputs, like in the case in Bernoulli or Markov processes.
The introduction of other bioinspired learning algorithms, namely STDP and SDSP provides a similar level of accuracy, in a similar computational time, with the exception of the SDSP algorithm for Markov processes, where the computational time is an order of magnitude longer. However, the lower MSE, and MAE for Bernoulli and Markov processes indicate that STDP fine-tunes synaptic weights more effectively than Hebbian learning, reducing classification errors.
Unsupervised learning algorithms help to achieve high accuracies with much fewer neurons in layers for uncorrelated data.
TABLE II: Influence of unsupervised learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Unsupervised Hebbian Bernoulli Yes 10 128 51.93 99.00 0.0100 0.0100 0.9600
Markov Yes 10 128 74.66 95.78 0.0422 0.0422 0.8313
Poisson Yes 10 64 17.17 90.00 0.1000 0.1000 0.5998
STDP Bernoulli Yes 10 64 62.08 99.50 0.0050 0.0050 0.9800
Markov Yes 10 128 513.73 98.00 0.0200 0.0200 0.9200
Poisson Yes 10 64 63.36 89.50 0.1050 0.1050 0.5798
SDSP Bernoulli Yes 10 128 43.2 98.00 0.0200 0.2000 0.9200
Markov Yes 10 128 46.1 93.87 0.0613 0.0613 0.7549
Poisson Yes 10 64 17.83 92.50 0.0750 0.0750 0.6999
Table III presents the influence of the superseded learning algorithm such as backpropagation, tempotron, SpikeProp, Chronotron and ReSuMe on the performance of SNNs. The BP algorithm achieves perfect classification accuracy for Bernoulli, Markov, and Poisson processes, as evidenced by zero MSE and MAE, indicating no classification errors. The $R^{2}$ score of 1.000 further confirms its ability to perfectly predict sequence outputs. However, this unmatched accuracy comes at a significant computational cost, with processing times, making BP the most resource-intensive supervised method.
Although BP is theoretically optimal in terms of predictive performance, its high computational demands render it impractical for real-time or large-scale applications where efficiency is crucial. This trend also holds for the STBP algorithm, but with a significantly greater time expenditure than BP. In turn, in [11], the authors state that they studied Bernoulli sources; however, in their network, the input layer consisted of LIF neurons, while the remaining layers comprised perceptrons trained using the tempotron algorithm. They did not present results for this datasets. We achive the high accuracy of clasiffication, reaching 99.00% for Bernoulli process and 100.00% for Markov process, while in the case of Poisson process hte level of accuracy in lower, namely 97.50%. In the case of tempotron, for Bernoulli and Poisson the calculations are very fast, taking about 7s. For Markov the time is eight times longer, while classification depends not just on the current state but also on prior states, the model must process sequences over multiple time steps, significantly increasing computational demands. Since the Poisson process does not have explicit dependencies like Markov sequences, computations remain efficient.
Also, the computation for the tempotron also requires fewer layers of neurons than BP and STBP with a slight loss of precision. In turn, the SpikeProp algorithm requires fewer neurons in the layers, which affects the computation speed for Markov and Poisson processes. This number decreases for all datasets when we apply the Chronotron learning algorithm, but the computation time is longer only in the case of the Poisson algorithm. Moreover, unlike classical backpropagation, which requires full gradient updates over all layers, Spikprop optimizes synaptic weights through event-driven updates, significantly reducing computational complexity in comparison to BP and STBP.
The ReSuMe demontrate the lowest average classification accuracy among all evaluated supervised learning algotihms. Is quite computationally efficient, while relies on spike pattern association through supervised reinforcement. Consequently, this results in less precise weight corrections, leading to higher classification error.
Supervised biologically inspired learning algorithms help achieve high accuracies with much fewer neurons in layers for uncorrelated data than supervised algorithms, non-biologically inspired learning algorithms, and unsupervised algorithms.
TABLE III: Influence of supervised learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Supervised BP Bernoulli No 10 128 866.44 99.00 0.0100 0.0100 0.9600
Markov No 10 256 6211.62 100.00 0.0000 0.0000 1.0000
Poisson No 10 512 12916.37 97.50 0.0250 0.0250 0.9000
STBP Bernoulli No 10 128 2261.43 100.00 0.0000 0.0000 1.0000
Markov No 10 512 48258.21 100.00 0.0000 0.0000 1.0000
Poisson No 10 512 30093.49 96.50 0.0350 0.0350 0.8599
tempotron Bernoulli Yes 10 128 7.0 96.00 0.0400 0.0400 0.8399
Markov Yes 10 128 40.5 90.59 0.0941 0.0941 0.6234
Poisson Yes 10 128 7.23 89.50 0.1050 0.1050 0.5798
Spikeprop Bernoulli Yes 10 128 64.28 100.00 0.0000 0.0000 1.000
Markov Yes 10 64 47.0 92.34 0.0766 0.0766 0.6935
Poisson Yes 10 32 20.15 91.50 0.0850 0.0850 0.6599
Chronotron Bernoulli Yes 10 32 18.61 98.50 0.01500 0.01500 0.9400
Markov Yes 10 32 46.30 92.34 0.9300 0.0175 0.9300
Poisson Yes 10 32 113.98 90.50 0.0950 0.0950 0.6198
ReSuMe Bernoulli Yes 10 128 68.55 92.50 0.0750 0.0750 0.6999
Markov Yes 10 64 46.05 91.60 0.0840 0.0840 0.6642
Poisson Yes 10 64 21.45 91.50 0.0850 0.0850 0.6590
Table IV presents the influence of hybrid learning algorithms such as ANN-SNN conversion, STPD based on rewards, and BAL on SNN performance. For ANN-SNN conversion LZC classifier achieves high classification accuracy for all datasets. However, for Poisson process has slightly lower accuracy due to its high randomness. It is worth to stress that the Markov process requires the longest processing time due to its sequential dependencies. Moreover, the Poisson process is processed fast, despite being more stochastic. Additionally, the higher errors for Poisson data are secondary to the fact that randomness complicates the model is introduced at the boundary limit.
We achieve comparable results with the reward-based and BAL learning algorithms. However in the case of reward based, for Markov and Poisson processes the network requires more neurons in layers than ANN-SNN conversion, but the computation time is shorter. For Bernoulli the network requires twice as few neurons in layers, but the computation time is five times shorter. BAL has six times longer computation time for a Poisson process than the reward based algorithm. Unlike standard Hebbian learning or STDP, BAL seems to deal with stochastic volatility more efficiently.
Thus, Poisson-distributed data presents the greatest challenge but also offers the most biologically realistic testing ground. Across nearly all algorithms, performance was lower for Poisson inputs compared to Bernoulli or Markov, indicating difficulties in processing highly irregular spike trains. It is worth highlighting that biologically inspired and hybrid algorithms like SDSP and BAL demonstrated comparatively better adaptability, underscoring their value for neuromorphic systems. Additionally, the computational burden for Poisson inputs was higher for gradient-based approaches due to lack of structure, which complicates error propagation and increases time complexity.
TABLE IV: Influence of hybrid learning algorithm on SNN.
Type Subtype Datasets Bio-inspired Epochs $n$ Time [s] Accuracy [%] MSE MAE $R^{2}$ score
Hybrid ANN-SNN Bernoulli No 10 64 21.41 99.50 0.0050 0.0050 0.9800
Markov No 10 32 37.25 92.58 0.0742 0.0724 0.7031
Poisson No 10 32 115.41 94.50 0.0550 0.0550 0.7799
Reward Bernoulli No 10 32 4.3 99.50 0.0050 0.0050 0.9800
Markov No 10 64 36.32 90.35 0.0965 0.0965 0.9948
Poisson No 10 64 14.83 91.00 0.0900 0.0900 0.6399
BAL Bernoulli Yes 10 32 1.56 97.00 0.0300 0.0300 0.8800
Markov Yes 10 128 46.37 95.38 0.0462 0.0462 0.8153
Poisson Yes 10 128 95.98 94.50 0.0550 0.0550 0.7799
V Conclusions
This comprehensive benchmarking study sheds light on the capabilities, trade-offs, and biological plausibility of various learning strategies in Spiking Neural Networks. It reveals that no single algorithm universally outperforms others across all metrics, data types, or computational conditions.
Higher input correlations, as observed for Poisson-distributed data, tend to reduce classification accuracy, emphasizing the inherent difficulty of extracting informative features from less structured signals. In contrast, more structured input types such as Bernoulli and Markov sequences facilitate greater learning stability and improved performance.
Hebbian learning remains computationally efficient and accurate for structured data, but struggles with Poisson-distributed inputs. STDP and SDSP, grounded in spike timing, provide biologically plausible yet effective mechanisms for learning from both structured and moderately unstructured inputs. Algorithms like BP and STBP are unmatched in precision but are computationally expensive, particularly for inputs like Poisson sequences. Their reliance on full backpropagation limits real-time neuromorphic applicability, despite theoretical optimality. ANN-SNN conversion trough a LZC and reward-based STDP offer scalable and flexible solutions. BAL, in particular, balances biological realism and computational efficiency, emerging as a promising method for future biologically plausible SNNs. Poisson inputs offer a realistic benchmark but challenge traditional learning rules. Hybrid and bio-inspired algorithms like BAL and SDSP demonstrate strong adaptability to these conditions, offering a viable path toward more robust neuromorphic learning systems.
Flexible hybrid learning architectures that dynamically adapt learning strategies in response to input characteristics hold promise for enhancing the performance of spiking neural networks under diverse operational conditions. Advancing our understanding of spike contributions in non-differentiable learning contexts remains a key limitation. Integrating event-driven learning mechanisms with biologically inspired regularization may offer promising directions for enhancing optimization efficiency in SNNs. Moreover, real-time evaluation of the trade-offs between accuracy and computational efficiency is essential for effective integration of SNNs on neuromorphic hardware, particularly in edge and low-power computing environments. To approach functional parity with biological intelligence, future learning models must better emulate the noise robustness and generalization capabilities of the brain, in particular under stochastic, Poisson-like input regimes.